A manufacturer of Virtual Reality headsets has seen sales increase from 25,000 units per year to 100,000 units per year in 8 years. What has been the firm's average annual rate of increase in the number of television sets sold? Use the Rule of 72 to determine your answer.

The average annual rate of change has been between 10% and 11%.

The average annual rate of change has been between 18% and 19%.

The average annual rate of change has been between 15% and 16%.

There is not enough information to answer this question.
Answer #1

As sales has quadrupled in 8 years, that mean sales doubled in 4 years.

So using Rule of 72, rate of interest to double the return in t years = 72/t

So, r = 72/4 = 18%

So, option B is correct